OpenStudy (anonymous):

Translate the following phrases into algebraic expressions. Explain, in complete sentences, the difference between the two. (1.) 2 minus 3 times x (2.) 2 less than 3 times x

OpenStudy (jonnyvonny):

The first one is 2-3x. Second one is 3x-2. The first one is a sequence; it literally reads "2 - 3*x" The second one is retardedly worded, and basically means, "First, after you multiply 3 and x, you then take away 2".

OpenStudy (anonymous):

Thank you !

OpenStudy (jonnyvonny):

No problem, if you have any more questions, feel free to ask.

OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 2x − 3y − z if x = −2, y = 3, and z = −2 −11 11 −15 15

OpenStudy (jonnyvonny):

This you just literally plug in, and follow PEMDAS: 2(-2)-3(3)-(-2) -4-9+2=?